Flint Company’s record of transactions for the month of April was as follows.

Purchases

Sales

April 1 (balance on hand) 708 @ $5.00 April 3 590 @ $10.00
4 1,770 @ 5.09 9 1,652 @ 10.00
8 944 @ 5.38 11 708 @ 11.00
13 1,416 @ 5.48 23 1,416 @ 11.00
21 826 @ 5.68 27 1,062 @ 12.00
29 590 @ 5.87 5,428
6,254

Assuming that periodic inventory records are kept in units only, compute the inventory at April 30 using LIFO and average-cost. (Round answer to 0 decimal places, e.g. 2,760.)

LIFO $

Entry field with incorrect answer?

Average-cost $

Entry field with correct answer4427
Answer #1

Closing inventory units = 6,254 - 5,428 = 826.

LIFO Method - Periodic Inventory

Closing inventory = ( 708 * $ 5 ) + ( 118 * $ 5.09 ) = $ 4,141.

Average cost Method.

Cost of goods available for sale = $ 33543

Average cost = $ 33543 / 6,254 = $ 5.4

Closing Inventory = 826 * $ 5.4 = $ 4461